# Threshold Filtering Packing for Supervised Fine-Tuning: Training Related Samples within Packs

### v1

@@ 5.1 TFP Design
We conducted a series of ablation studies on several critical design choices for TFP using the GSM8K dataset due to its widespread use and high quality. Embedding models play a pivotal role in these ablation studies, as they capture different semantic meanings, which is essential for understanding the operational mechanisms of TFP.
Initially, we evaluated various embedding models, including bert-base-uncased Devlin et al. (2019), RoBERTa Liu et al. (2019), all-MiniLM-L6-v2, E5-mistral-7b-instruct Wang et al. (2022), and the hidden layers of Llama3 AI@Meta (2024). These models fall into two primary categories Liu et al. (2024): the Model-based method, which includes base models like bert-base-uncased, Llama3, and RoBERTa, and the Semantic-based method, which employs advanced sentence embedding models such as E5-mistral-7b-instruct and all-MiniLM-L6-v2.
As presented in Figure 4, the base models generally outperformed the others, with bert-base-uncased emerging as the most effective. This result suggests that base models are better suited for sample aggregation in contextual training, likely because TFP requires models that provide contextual information rather than those focused solely on semantic similarity as in clustering or retrieval tasks.
Further analysis explored the optimal data segments for TFP within the Alpaca format, which typically includes instruction, input, and output components. Given that inputs often lack substantial content, we conducted experiments on the instruction part, output part, and entire data. Our experiments on the GSM8K dataset indicated that using instructions for similarity calculations was the most beneficial. This finding is significant because calculating similarity within the instruction section reduces the risk of cross-contamination. By avoiding the clustering of similar outputs, we can prevent models from simply reproducing previous outputs since SFT only calculates loss on the outputs.
@@ 5.2 The distance threshold number $r$
We now examine the influence of the distance threshold number $r$ on the TFP algorithm. The parameter analysis were conducted using Llama3-8B and Llama2-7B to determine how variations in $r$ impact performance within models from the same series. The results, evaluated on the GSM8K dataset, are presented in Figure 4.
Focusing solely on $r$, setting $r=0$ causes a greedy selection of the nearest samples (equivalent to TSP), which leads to reduced performance. As $r$ increases, performance tends to degrade, eventually resembling that of random packing. For different models, while a smaller $r$ may benefit Llama2-7B by maintaining contextual relevance, Llama3-8B may require a larger $r$ to avoid the pitfalls of excessive similarity. This may occur because a more powerful model, being more sensitive to contextual reasoning, is more prone to performance degradation when exposed to overly similar or repetitive data.
@@ 6 Conclusion
We present TFP, a novel approach for packing samples during the SFT phase. This method exposes language models to relevant samples, allowing them to learn from relevant context and effectively adapt to few-shot evaluation. TFP is highly scalable and simple-to-implement, compatible with any existing SFT framework by simply altering the sequence of samples in the packing process. Our comprehensive evaluation demonstrates that TFP significantly enhances SFT performance for both text and tabular data, particularly in few-shot tasks. Additionally, TFP improves fairness in predictions. By adjusting the ratio within packs, TFP introduces samples with similar texts but varying sensitive attributes, which helps mitigate biases in LLMs without sacrificing accuracy.
@@
Moreover, we have not fully explored the relationship between providing related context within packs and other stages of training. Recent studies highlight the importance of maintaining internal knowledge consistency before and after SFT Ren et al. (2024); Yang et al. (2024). Earlier research on pretraining language models with related documents has shown promising results Staniszewski et al. (2024); Shi et al. (2024); Yasunaga et al. (2022); Yu et al. (2022). These methods either use metadata or retrieval techniques to group mutually relevant documents into long, coherent training examples. Our experiments indicate that providing context during SFT stages enhances in-context learning. We plan to explore integrating TFP with pretraining and in-context learning methods in future work.
@@ Appendix B Experiment Details
We use the hyperparameters from the original framework for SFT, and the parameters of TFP are set in Section 5.
@@ Appendix C Evaluation Details
    You are a helpful, pattern-following assistant. For each snippet of text, predict whether the text of the individual comment is harmless or toxic. The answer should be exactly ’harmless’ or ’toxic’." + input_text + "Predict whether the above comment is harmless or toxic:
@@
    A person in 1996 has the following attributes: age 29, workclass Private, final weight 85572, education Bachelors, highest education level 13, marital status Married-civ-spouse, occupation Exec-managerial, relationship Wife, race White, sex Female, capital gain 0, capital loss 0, hours per week 42, native country United-States.predict whether the annual income of the person is greater than \$50k:

### v2

@@ 5.1 TFP Design
We conduct a series of ablation studies on several critical design choices for TFP using the GSM8K dataset due to its widespread use and high quality. Embedding models play a pivotal role in these ablation studies, as they capture different semantic meanings, which is essential for understanding the operational mechanisms of TFP.
Initially, we evaluate various embedding models, including bert-base-uncased Devlin et al. (2019), RoBERTa Liu et al. (2019), all-MiniLM-L6-v2, E5-mistral-7b-instruct Wang et al. (2022), and the hidden layers of Llama3 AI@Meta (2024). These models can be divided into two main approaches Liu et al. (2023): Model-based methods, such as bert-base-uncased, Llama3, and RoBERTa, and Semantic-based methods, like E5-mistral-7b-instruct and all-MiniLM-L6-v2.
As shown in Figure 4, the model-based approaches generally outperform the semantic-based ones, with bert-base-uncased achieving the best results. This suggests that base models are more effective for sample aggregation in contextual training, likely because TFP relies on models that capture contextual information. In contrast, models focused primarily on semantic similarity, like those used in clustering or retrieval tasks, do not perform well.
Further analysis explores the optimal data segments for TFP within the Alpaca format, which typically includes instruction, input, and output components. Given that inputs often lack substantial content, we conduct experiments on the instruction part, output part, and entire data. Our experiments on the GSM8K dataset indicate that using instructions for similarity calculations is the most beneficial. This finding is significant because calculating similarity within the instruction section reduces the risk of cross-contamination. By avoiding the clustering of similar outputs, we can prevent models from simply reproducing previous outputs since SFT only calculates loss on the outputs.
@@ 5.2 The distance threshold number $r$
We analyze the effect of the distance threshold number $r$ on the TFP algorithm. The parameter analysis were conducted using Llama3-8B and Llama2-7B to determine how variations in $r$ impact performance within models from the same series. The results, evaluated on the GSM8K dataset, are presented in Figure 4.
Setting $r=0$ results in a greedy selection (equivalent to TSP), showing degraded performance. As $r$ becomes too large, performance declines further, approaching random packing. An appropriate $r$ helps Llama2-7B and Llama3-8B to maintain contextual relevance and avoid excessive similarity. Llama3-8B needs a larger $r$ than Llama2-7B because it is a more powerful model and, therefore, more sensitive to contextual reasoning and more prone to performance degradation when exposed to overly similar or repetitive data.
@@ 6 Conclusion
We present TFP, a novel and scalable method for packing samples during SFT, enabling language models to learn from relevant context and adapt effectively to few-shot evaluation. TFP is simple-to-implement and integrates with existing SFT framework by adjusting the sample sequence. Our evaluation shows TFP significantly boosts SFT performance on both text and tabular data, especially in few-shot tasks. It also improves fairness by introducing similar texts with varying sensitive attributes, helping reduce biases in LLMs without compromising accuracy.
@@
Moreover, we have not fully explored the relationship between providing related context within packs and other stages of training. Recent studies highlight the importance of maintaining internal knowledge consistency before and after SFT Ren et al. (2024); Yang et al. (2024). Earlier research on pretraining language models with related documents has shown promising results Staniszewski et al. (2023); Shi et al. (2023); Yasunaga et al. (2022); Yu et al. (2022); Zhao et al. (2024b). These methods either use metadata or retrieval techniques to group mutually relevant documents into long, coherent training examples. Our experiments indicate that providing context during SFT stages enhances in-context learning. We plan to explore integrating TFP with pretraining and in-context learning methods in future work.
@@
Acknowledgments
This work is supported by the National Science Foundation (NSF) Grant #2312862, NSF-Simons SkAI Institute, National Institutes of Health (NIH) #R01AG091762, and a Cisco gift grant.
@@ Appendix B Experiment Details
We use the hyperparameters from the original framework for SFT, and the parameters of TFP are set in Section 5. The threshold value corresponds to the nearest 2% of all pairwise distances, ensuring that TFP-generated packs have significantly reduced average distances compared to random packing and the greedy TSP algorithm.
@@ Appendix C Evaluation Details
    You are a helpful, pattern-following assistant. For each snippet of text, predict whether the text of the individual comment is harmless or toxic. The answer should be exactly ’harmless’ or ’toxic’. Predict whether the comment is harmless or toxic:
@@
    Predict whether the annual income of the person is greater than \$50k:
Appendix D Analysis of Packs
This appendix provides further details regarding the methods used to generate data packs for SFT.
D.1 Comparison of Packing Methods
We compared different packing methods by calculating the average distances between data points within each pack using bert-base-uncased as the embedding model. The results for GSM8K are summarized in the Table 5.
Packing Method Average Distance
Random Order 9.23
TSP Order 6.18
TFP Order 6.48
$k$-NN (limiting overlap) 8.07
Table 5: Average distances between data points within each pack for different packing methods on the GSM8K dataset.
The results show that TFP outperforms other methods in preserving meaningful context, as indicated by the lower average distance. The $k$-NN method, even with limiting overlap, results in more dispersed packs, making it less effective for fine-tuning tasks.
D.2 Graph Creation
A fully connected graph is used for packing, where each node represents a data sample, and edges are weighted based on pairwise cosine similarity between sample embeddings. This graph construction ensures robust and diverse pack generation.
Initial node selection has minimal impact on TFP’s performance due to the diversity-focused packing strategy. Additionally, the randomized training order of packs further minimizes the effect of the starting point.
Appendix E Other Results
To evaluate TFP’s generalizability, we conducted experiments on Llama3-8B using two additional training datasets: MathInstruct and CodeAlpaca-20k.
Method MATH GSM8K
Vanilla FT 20.5 63.9
Random packing 18.4 56.6
TFP 22.7 68.2
Table 6: Results on MathInstruct datasets.
Method MBPP HumanEval
Vanilla FT 45.4 38.4
Random packing 45.4 39.6
TFP 49.4 42.1
Table 7: Results on CodeAlpaca-20k datasets.
As shown, TFP consistently outperforms other methods across all benchmarks, demonstrating its adaptability to various datasets.